# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ Appendix A Derivations for Table 1: Spikiness, conditional perturbation rate, and quantile growth
Consequently, the conditional perturbation modulus for the mixture satisfies
$L_{CP}(M)\ \leq\ \max_{k}M_{\eta,k}(M)\ +\ \frac{1}{\pi_{\min}}\sup_{k}\|\nabla_{\eta}\pi_{k}\|_{L^{\infty}(K_{\eta})}.$
The last term is a constant, giving the result.
Appendix B Proofs
In this Section we provide proofs of the general results on conditional normalizing flows and NeVI-Cut solutions using them. Proofs of the results for specific flow models (Theorems 2 and 3) are in the next Section.
@@ Appendix B Proofs
Lipschitz modulus for $\log\partial_{z}T_{p}$. For $(\eta_{i},z_{i})\in K_{M}$ set $\theta_{i}=T_{p}(\eta_{i},z_{i})$. Applying Lemma 2, by the two moduli condition on $K_{\eta}\times[-G(M),G(M)]$ and since $\log\phi$ is $N$ Lipschitz on [M,M][-M,M],
@@
By the mean value theorem and Definition 1(iii)(b), $|\theta_{1}-\theta_{2}|=\big|\int_{z_{1}}^{z_{2}}\partial_{u}T_{p}(\eta_{1},u)\,du\big|\leq e^{H(M)}|z_{1}-z_{2}|$. Thus
@@
Therefore $\log\partial_{z}T_{p}$ is $\mathsf{L}_{b}(M)$ Lipschitz on $K_{M}$ with respect to the $\ell_{1}$ distance $\|(\eta,z)-(\eta^{\prime},z^{\prime})\|_{1}:=\|\eta-\eta^{\prime}\|_{1}+|z-z^{\prime}|$.
Lipschitz modulus for $T_{p}(\eta,0)$ in $\eta$. FOr simplicity of notaiton, we use $F_{\eta}(\cdot)$ to denote the conditional cdf $F_{p}(\cdot\,|\,\eta)$. Fix $\eta_{1},\eta_{2}\in K_{\eta}$ and set $u=\Phi(0)$, $\theta_{i}=T(\eta_{i},0)=F_{\eta_{i}}^{-1}(u)$, and $g=F_{\eta_{2}}^{-1}$. Then
@@
Combining the last two displays gives
@@
Proof of Lemma 2.
We have $F_{p}\!\big(T_{p}(\eta,z)\,|\,\eta\big)=\Phi(z).$ Differentiate both sides with respect to $z$ and use the chain rule:
$\partial_{\theta}F_{p}\!\big(T(\eta,z)\,|\,\eta\big)\cdot\partial_{z}T_{p}(\eta,z)=\phi(z).$
$p\!\big(T(\eta,z)\mid\eta\big)\,\partial_{z}T(\eta,z)=\phi(z).$
@@
$d_{M}(p,q):=\sup_{(\eta,z)\in K_{M}}\max\Big(\,|T_{p}(\eta,0)-T_{q}(\eta,0)|,\;|\log\partial_{z}T_{p}(\eta,z)-\log\partial_{z}T_{q}(\eta,z)|\,\Big).$
Step 1: Core uniform approximation on $K_{M}$. By Assumption 2(iii), for budget $m=C(M,\varepsilon)$ there exists $T\in\mathcal{F}_{m}$ such that
$d_{M}\big(q_{T},p\big)\ =\ \sup_{(\eta,z)\in K_{M}}\max\!\Big(\,|T(\eta,0)-T_{p}(\eta,0)|,\;|\log\partial_{z}T(\eta,z)-\log\partial_{z}T_{p}(\eta,z)|\,\Big)\ \leq\ \varepsilon.$ (21)
Write $\Delta_{0}(\eta):=T(\eta,0)-T_{p}(\eta,0),\qquad\Delta_{\ell}(\eta,z):=\log\partial_{z}T(\eta,z)-\log\partial_{z}T_{P}(\eta,z).$ Then $|\Delta_{0}|\leq\varepsilon$ and $|\Delta_{\ell}|\leq\varepsilon$ on $K_{M}$.
@@
On $|Z|\leq M$, the first difference is $\leq\varepsilon$ by (21). For the second, use the spikiness modulus on the bounded $\theta$–range: on $|z|\leq M$,
@@
because $|\log\partial_{u}T_{p}|\leq H(M)$ and $|e^{\Delta_{\ell}}-1|\leq e^{\varepsilon}-1\leq(e-1)\varepsilon\leq 2\varepsilon$ for $\varepsilon\leq 1/2$. Note that on $K_{M}$, $|T_{p}(\eta,Z)|\leq G(M):=M_{0}+|M|e^{A_{d}+B_{d}|M|}$ by (16). Similarly, $|T(\eta,Z)|\leq\widetilde{G}(M):=M^{*}+|M|e^{A^{*}+B^{*}|M|}$. As $M^{*}>M_{0}$, $A^{*}>A_{d}$, and $B^{*}>B_{d}+C_{d}$, we have $\widetilde{G}(M)\geq G(M)$ and both $T_{p}(\eta,Z)$ and $T(\eta,Z)$ lie in $\widetilde{G}(M),\widetilde{G}(M)]$. Hence,
$\big|\log p\!\big(T(\eta,z)\mid\eta\big)-\log p\!\big(T_{p}(\eta,z)\mid\eta\big)\big|\leq L_{S}\!\big(\widetilde{G}(M)\big)\,|T(\eta,z)-T_{p}(\eta,z)|\leq\varepsilon\,L_{S}\!\big(\widetilde{G}(M)\big)\,(1+2Me^{H(M)}).$
@@
$\leq\varepsilon\Big\{1+L_{S}\!\big(\widetilde{G}(M)\big)(1+2Me^{H(M)})\Big\}$
@@
Step 3: Gaussian tails $|z|>M$. Using the identity above and Assumption 2(ii),
$\log\partial_{z}T(\eta,z)\leq A^{*}+B^{*}|z|,\qquad|T(\eta,z)|\leq|T(\eta,0)|+\!\int_{0}^{|z|}\!\partial_{u}T(\eta,u)\,du\leq M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}.$
This leads to
$T(\eta,z)^{2}$ $\;\leq\;\Big(M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}\Big)^{2}$
 $\;\leq 2M^{*2}+2e^{2A^{*}}z^{2}e^{2B^{*}|z|}.$
By the tail–thinness lower bound in Definition 1(iv), $\log p\!\big(T(\eta,z)\mid\eta\big)\leq\alpha\,T(\eta,z)^{2}-\log c$. With $\log\phi(z)=-\tfrac{z^{2}}{2}-\tfrac{1}{2}\log(2\pi)$, Definition 1(iv) and Assumption 2(ii) give, for $|z|>M$ and all $\eta$,
$\log\phi(z)-\log p\!\big(T(\eta,z)\mid\eta\big)-\log\partial_{z}T(\eta,z)$ $\leq-\frac{z^{2}}{2}-\frac{1}{2}\log(2\pi)-\log c+\alpha\,T(\eta,z)^{2}+A^{*}+B^{*}|z|$
@@
$C_{1}=A^{*}-\tfrac{1}{2}\log(2\pi)-\log c+2\alpha\,M^{*2},\qquad C_{2}=\alpha 2e^{2A^{*}}.\;$
@@
$\int_{|z|>M}\!\phi(z)\,[\cdots]\,dz$ $\leq\frac{1}{\sqrt{2\pi}}\Big\{C_{1}\,I_{0}(M)\;+\;C_{2}\,I_{2}(M)\;+\;B^{*}\,I_{1}(M)\Big\},$
@@
$I_{0}(M)\leq\frac{2\,\phi(M)}{M},\qquad I_{1}(M)=2\!\int_{M}^{\infty}\!ze^{-z^{2}/2}dz=2\,e^{-M^{2}/2}.$
For $I_{2}$, Using $(t+\gamma)^{2}\leq 2(t^{2}+\gamma^{2})$, we have
$1+(t+\gamma)^{2}\;\leq\;1+2t^{2}+2\gamma^{2}.$
Hence
$I_{2}(M)=2e^{\gamma^{2}/2}\int_{M-\gamma}^{\infty}\!\bigl(1+(t+\gamma)^{2}\bigr)e^{-t^{2}/2}\,dt\;\leq\;2e^{\gamma^{2}/2}M_{\gamma}\!\int_{M-\gamma}^{\infty}\!(1+t^{2})e^{-t^{2}/2}\,dt,$
where $M_{\gamma}:=\max\{1+2\gamma^{2},\,2\}$.
@@
with $C_{\gamma}:=4M_{\gamma}e^{\gamma^{2}/2}$.
Setting $\gamma=2B^{*}$ and $M_{\gamma}\leq 1+2\gamma^{2}+2\leq 3+8B^{*2}$, yields
$I_{2}(M)\;\leq\;\widetilde{C}_{B^{*}}\,(1+M)\,e^{-\frac{1}{2}\,(M-2B^{*})^{2}},\mbox{ where }\widetilde{C}_{B^{*}}=4(3+8B^{*2})e^{2B^{*2}}.$
@@
Absorbing $2e^{1/2}(1+2B^{*})$ in $R$ we can choose $M=2B^{*}+1+\sqrt{\,2\log\!\frac{R}{\epsilon}}$ for a suitable constant $R$. So, for any small $\varepsilon^{\prime}$, $M$ can be expressed as $O(\sqrt{-\log\varepsilon^{\prime}})$ where the constant depends only on $M^{*},A^{*},B^{*},\alpha$ and $c$. ∎
@@
(a) Existence. Recall that the parameter set $\Theta_{\,m}\subset\mathbb{R}^{m}$ is compact by Assumption 2. For each fixed $(\eta,z)$, the maps $\vartheta\mapsto T_{\vartheta}(\eta,z)$ and $\vartheta\mapsto\partial_{z}T_{\vartheta}(\eta,z)$ are continuous by (Assumption 2(i)), hence the integrand
@@
is continuous in $\vartheta$. By Assumption 2(ii) and Definition 1(iv),
@@
with $C_{1},C_{2}$ independent of $\vartheta$. Since $\int_{\mathbb{R}}(1+z^{2})e^{2B^{*}|z|}\phi(z)\,dz<\infty$ (complete the square), the envelope is integrable against $\phi(z)p(\eta)$. By dominated convergence, $\vartheta\mapsto R_{N}(q_{T_{\vartheta}})$ is continuous on the compact set $\Theta_{\,m}$, hence it attains its minimum; existence of $\hat{q}_{m}$ follows.
(b) KLD bound. By Theorem 1, given a small $\varepsilon$, for the budget $m:=C^{*}(\varepsilon)$ there exists $T\in\mathcal{F}_{m}$ such that $R_{m}\big(q_{T}\big)\;\leq\;\varepsilon.$ Because $\hat{q}_{m}$ minimizes $R(q)$ over $q\in\mathcal{Q}_{m}$,
@@
Since $\mathcal{Q}_{m}=\{q_{T_{\vartheta}}:\vartheta\in\Theta_{m}\}$, with a slight abuse of notation, we can write $R(q)$ as $R(\vartheta)$ and $R_{N}(q)$ as $R_{N}(\vartheta)$ for the corresponding $\vartheta$. As $\Theta_{m}$ is compact and $\vartheta\mapsto R_{N}(q_{\vartheta})$ is continuous, by Weierstrass Extreme Value Theorem, we have $\hat{q}_{N}\in\arg\min_{\mathcal{Q}_{m}}R_{N}(q)$.
@@
Let $p(\theta\,|\,\eta)$ satisfies Assumption 1 for some compact $K_{\eta}$ and $p(\eta)$ is a distribution supported on $K_{\eta}$. Let $\{\eta^{(r)}\}_{r\geq 1}$ be a stationary ergodic Markov chain with invariant distribution $p(\eta)$. Let $\mathcal{F}_{m}=\{T_{\vartheta}|\;\vartheta\in\Theta_{m}\}$ denote a class satisfying Assumption 2. Let $q_{\vartheta}$ to be the conditional distribution $\theta\,|\,\eta=T_{\vartheta}(\eta,Z),Z\sim N(0,1)$ and
$f_{\vartheta}(\eta)=D_{KL}\!\big(q_{\vartheta}(\theta\mid\eta)\,\|\,p(\theta\mid\eta)\big),\qquad R(\vartheta)=\mathbb{E}[f_{\vartheta}(\eta)],\quad R_{N}(\vartheta)=\frac{1}{N}\sum_{r=1}^{N}f_{\vartheta}(\eta^{(r)}).$
@@
Envelope. Let $H(\vartheta,\eta,z)=\log\phi(z)-\log p\!\big(T(\eta,z)\mid\eta\big)-\log\partial_{z}T(\eta,z)$ and $f_{\vartheta}(\eta)=\mathbb{E}[H(u,\eta,Z)]$. Following the proof of Corollary 1, $f_{\vartheta}(\eta)\leq\tau$ for some positive $\tau$ free of $\eta$ and for each fixed $\eta$, the map $\vartheta\mapsto f_{\vartheta}(\eta)$ and is continuous on compact $\Theta_{m}$. Define for $\varepsilon>0$
@@
Then $L_{\varepsilon}(\eta)\to 0$ pointwise as $\varepsilon\downarrow 0$, and $0\leq L_{\varepsilon}(\eta)\leq 2\tau$.
@@
$\mathbb{E}\,L_{\varepsilon}(\eta)\longrightarrow 0\qquad(\varepsilon\downarrow 0).$ (22)
Since $\{\eta^{(r)}\}_{r\geq 1}$ is stationary ergodic and $L_{\varepsilon}\in L^{1}(p)$, Birkhoff’s ergodic theorem yields
$\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta^{(r)})\;\xrightarrow[N\to\infty]{a.s.}\;\mathbb{E}_{p(\eta)}[L_{\varepsilon}(\eta)].$
@@
$\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\big|R_{N}(\vartheta)-R_{N}(\vartheta_{j})\big|+\big|R(\vartheta)-R(\vartheta_{j})\big|.$
$\leq\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta^{(r)})+\mathbb{E}\,L_{\varepsilon}(\eta).$
@@
$\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\max_{1\leq j\leq n(\varepsilon)}\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta^{(r)})+\mathbb{E}\,L_{\varepsilon}(\eta).$
Taking $\limsup_{N\to\infty}$ on both sides and applying the ergodic theorem to the empirical average of $L_{\varepsilon}(\eta^{(r)})$ yields
$\limsup_{N\to\infty}\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\max_{1\leq j\leq n(\varepsilon)}\limsup_{N\to\infty}\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+2\,\mathbb{E}\,L_{\varepsilon}(\eta).$ (23)
For each fixed $j$, the ergodic theorem gives $R_{N}(\vartheta_{j})\to R(\vartheta_{j})$ almost surely, hence the first term on the right side of (23) is zero. Therefore
@@
Let $q_{m}^{\star}\in\arg\min_{\mathcal{Q}_{m}}R(q)$, which exists by Corollary 1. Then, for all $M$,
$R(\hat{q}_{N,m})$ $\leq R_{N}(\hat{q}_{N,m})+\Delta_{N,m}\leq R_{N}(q_{m}^{\star})+\Delta_{N,m}\leq R(q_{m}^{\star})+2\Delta_{N,m}.$

### v2

@@ Acknowledgements
Hence, this term is uniformly bounded.
Appendix S4 Proofs
In this Section, we provide proofs of the general results on conditional normalizing flows and NeVI-Cut solutions using them. Proofs of the results for specific flow models (Theorems 2 and 3) are in the next Section. We start with proof of Lemma 2, as the result is used in other proofs.
@@
By definition (12), $T_{p}(\eta,z)$ is a composition of $u\mapsto F_{p}^{-1}(u\mid\eta)$ and $z\mapsto\Phi(z)$. Under Assumption 1, $T_{p}$ is continuous on $K_{\eta}\times\mathbb{R}$, strictly increasing and continuously differentiable in $z$.
Following (12), we have $F_{p}\big(T_{p}(\eta,z)\,|\,\eta\big)=\Phi(z).$ Differentiate both sides with respect to $z$ and use the chain rule:
$\partial_{\theta}F_{p}\big(T_{p}(\eta,z)\,|\,\eta\big)\cdot\partial_{z}T_{p}(\eta,z)=\phi(z).$
$p\big(T_{p}(\eta,z)\mid\eta\big)\,\partial_{z}T_{p}(\eta,z)=\phi(z).$
Proof of Lemma 1.
Lipschitz modulus for $\log\partial_{z}T_{p}$. For $(z_{i},\eta_{i})\in K_{M}$ set $\theta_{i}=T_{p}(\eta_{i},z_{i})$. Applying Lemma 2, by the two moduli condition on $K_{\eta}\times[-G(M),G(M)]$ and since $\log\phi$ is $M$-Lipschitz on [M,M][-M,M],
@@
By the mean value theorem and Assumption 1(iii)(b)(iii)(b), $|\theta_{1}-\theta_{2}|=\big|\int_{z_{1}}^{z_{2}}\partial_{u}T_{p}(\eta_{1},u)\,du\big|\leq e^{H(M)}|z_{1}-z_{2}|$. Thus
@@
Therefore $\log\partial_{z}T_{p}$ is $\mathsf{L}_{b}(M)$-Lipschitz on $K_{M}$ with respect to the $\ell_{1}$ distance $\|(\eta,z)-(\eta^{\prime},z^{\prime})\|_{1}:=\|\eta-\eta^{\prime}\|_{1}+|z-z^{\prime}|$.
Lipschitz modulus for $T_{p}(\eta,0)$ in $\eta$. For simplicity of notation, we use $F_{\eta}(\cdot)$ to denote the conditional cdf $F_{p}(\cdot\,|\,\eta)$. Fix $\eta_{1},\eta_{2}\in K_{\eta}$ and set $u=\Phi(0)$, $\theta_{i}=T(\eta_{i},0)=F_{\eta_{i}}^{-1}(u)$, and $g=F_{\eta_{2}}^{-1}$. Then
@@
Combining the last two displays with the definition of $m_{0}$ gives
@@
$d_{M}(p,q):=\sup_{(z,\eta)\in K_{M}}\max\Big(\,|T_{p}(\eta,0)-T(\eta,0)|,\;|\log\partial_{z}T_{p}(\eta,z)-\log\partial_{z}T(\eta,z)|\,\Big).$
Step 1: Core uniform approximation on $K_{M}$. By Assumption 2(iii)(iii), for budget $m=C(M,\varepsilon)$ there exists $T\in\mathcal{F}_{m}$ such that
$d_{M}\big(q_{T},p\big)\ \ \leq\ \varepsilon.$ (S7)
Write $\Delta_{0}(\eta):=T(\eta,0)-T_{p}(\eta,0)$, and $\Delta_{\ell}(\eta,z):=\log\partial_{z}T(\eta,z)-\log\partial_{z}T_{p}(\eta,z)$. Then $|\Delta_{0}|\leq\varepsilon$ and $|\Delta_{\ell}|\leq\varepsilon$ on $K_{M}$.
@@
On $|Z|\leq M$, the first difference is $\leq\varepsilon$ by (S7). For the second, use the spikiness modulus on the bounded $\theta$–range: on $|z|\leq M$,
@@
because $|\log\partial_{u}T_{p}|\leq H(M)$ and $|e^{\Delta_{\ell}}-1|\leq e^{\varepsilon}-1\leq(e-1)\varepsilon\leq 2\varepsilon$ for $\varepsilon\leq 1/2$. Note that on $K_{M}$, $|T_{p}(\eta,Z)|\leq G(M):=M_{0}+|M|e^{A_{d}+B_{d}|M|}$ by (13). Similarly, $|T(\eta,Z)|\leq\widetilde{G}(M):=M^{*}+|M|e^{A^{*}+B^{*}|M|}$. As $M^{*}>M_{0}$, $A^{*}>A_{d}$, and $B^{*}>B_{d}$, we have $\widetilde{G}(M)\geq G(M)$ and both $T_{p}(\eta,Z)$ and $T(\eta,Z)$ lie in $\widetilde{G}(M),\widetilde{G}(M)]$. Hence,
$\big|\log p\big(T(\eta,z)\mid\eta\big)-\log p\big(T_{p}(\eta,z)\mid\eta\big)\big|$ $\leq L_{S}\big(\widetilde{G}(M)\big)\,|T(\eta,z)-T_{p}(\eta,z)|$
$\leq\varepsilon\,L_{S}\big(\widetilde{G}(M)\big)\,(1+2Me^{H(M)}).$
@@
$\leq\varepsilon\Big\{1+L_{S}\big(\widetilde{G}(M)\big)\left(1+2Me^{H(M)}\right)\Big\}$
@@
For this to be less than $\varepsilon^{\prime}/2$,
$\varepsilon\leq\frac{\varepsilon^{\prime}}{4\Big\{1+\widetilde{G}(M)L_{S}\big(\widetilde{G}(M)\big)\Big\}}.$
Step 3: Gaussian tails $|z|>M$. Using the identity above and Assumption 2(ii)(ii),
$\log\partial_{z}T(\eta,z)\leq A^{*}+B^{*}|z|,\quad|T(\eta,z)|\leq|T(\eta,0)|+\!\int_{0}^{|z|}\!\partial_{u}T(\eta,u)\,du\leq M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}.$
Using $(a+b)^{2}\leq 2(a^{2}+b^{2})$ for any real $a$ and $b$ leads to
$T(\eta,z)^{2}\;\leq\;\Big(M^{*}+e^{A^{*}}|z|e^{B^{*}|z|}\Big)^{2}\;\leq 2M^{*2}+2e^{2A^{*}}z^{2}e^{2B^{*}|z|}.$
By the tail-thinness lower bound in Assumption 1(iv)(iv),
$\log p\big(T(\eta,z)\mid\eta\big)\leq\alpha\,T(\eta,z)^{2}-\log c.$
With $\log\phi(z)=-\tfrac{z^{2}}{2}-\tfrac{1}{2}\log(2\pi)$, Assumption 1(iv)(iv) and 2(ii)(ii) give, for $|z|>M$ and all $\eta$,
$\log\phi(z)-\log p\big(T(\eta,z)\mid\eta\big)-\log\partial_{z}T(\eta,z)$
$\leq$ $-\frac{z^{2}}{2}-\frac{1}{2}\log(2\pi)-\log c+\alpha\,\left(2M^{*2}+2e^{2A^{*}}z^{2}e^{2B^{*}|z|}\right)+A^{*}+B^{*}|z|$
$\leq$ $-\frac{1}{2}\log(2\pi)-\log c+\alpha\,\left(2M^{*2}+2e^{2A^{*}}z^{2}e^{2B^{*}|z|}\right)+A^{*}+B^{*}|z|$
@@
$C_{1}=A^{*}-\tfrac{1}{2}\log(2\pi)-\log c+2\alpha\,M^{*2},\qquad C_{2}=2\alpha e^{2A^{*}}.\;$
@@
$\int_{|z|>M}\!\phi(z)\,[\cdots]\,dz$ $\leq\frac{1}{\sqrt{2\pi}}\Big[C_{1}\,I_{0}(M)\;+\;B^{*}\,I_{1}(M)\;+\;C_{2}\,I_{2}(M)\Big],$
@@
$I_{0}(M)\leq\frac{2\sqrt{2\pi}\,\phi(M)}{M},\qquad I_{1}(M)=2\!\int_{M}^{\infty}\!ze^{-z^{2}/2}dz=2\,e^{-M^{2}/2}.$
Rewriting $I_{2}$ and using $z^{2}-4B^{*}z=(z-2B^{*})^{2}-4{B^{*}}^{2}$,
$I_{2}(M)=2\int_{M}^{\infty}z^{2}e^{-z^{2}/2+2B^{*}z}dz=2e^{2{B^{*}}^{2}}\int_{M}^{\infty}z^{2}e^{-\frac{1}{2}(z-2B^{*})^{2}}dz.$
Using change of variable $t=z-\gamma$ where $\gamma=2B^{*}$, and $(t+\gamma)^{2}\leq 2(t^{2}+\gamma^{2})$, we have
$I_{2}(M)=2e^{\gamma^{2}/2}\int_{M-\gamma}^{\infty}(t+\gamma)^{2}e^{-t^{2}/2}\,dt\;\leq\;2e^{\gamma^{2}/2}M_{\gamma}\!\int_{M-\gamma}^{\infty}\!(1+t^{2})e^{-t^{2}/2}\,dt,$
where $M_{\gamma}:=\max\{2\gamma^{2},\,2\}$.
@@
with $C_{\gamma}:=4M_{\gamma}e^{\gamma^{2}/2}$. Setting $\gamma=2B^{*}$ and $M_{\gamma}\leq 2\gamma^{2}+2\leq 2(1+4B^{*2})$, yields
$I_{2}(M)\;\leq\;\widetilde{C}_{B^{*}}\,(1+M)\,e^{-\frac{1}{2}\,(M-2B^{*})^{2}},\mbox{ where }\widetilde{C}_{B^{*}}=8(1+4B^{*2})e^{2B^{*2}}.$
@@
Absorbing $2e^{1/2}(1+2B^{*})$ in $R$ we can choose $M=2B^{*}+1+\sqrt{\,2\log\!\frac{R}{\epsilon^{\prime}}}$ for a suitable constant $R$. So, for any small $\varepsilon^{\prime}$, $M$ can be expressed as $O(\sqrt{-\log\varepsilon^{\prime}})$ where the constant depends only on $M^{*},A^{*},B^{*},\alpha$ and $c$.
Combining the findings of steps 2 and 3 completes the proof.
@@
(a) Existence. Recall that the parameter set $\Theta_{\,m}\subset\mathbb{R}^{m}$ is compact by Assumption 2. For each fixed $(\eta,z)$, the maps $\vartheta\mapsto T_{\vartheta}(\eta,z)$ and $\vartheta\mapsto\partial_{z}T_{\vartheta}(\eta,z)$ are continuous by (Assumption 2(i)(i)), hence the integrand
@@
is continuous in $\vartheta$. By assumptions 2(ii)(ii) and 1(iv)(iv),
@@
with $C_{1},C_{2}$ independent of $\vartheta$. Since $\int_{\mathbb{R}}(1+z^{2})e^{2B^{*}|z|}\phi(z)\,dz<\infty$ (complete the square), the envelope is integrable against $\phi(z)p(\eta)$. By dominated convergence, $\vartheta\mapsto R(q_{T_{\vartheta}})$ is continuous on the compact set $\Theta_{\,m}$, hence by Weierstrass Extreme Value Theorem it attains its minimum; existence of $\hat{q}_{m}$ follows.
(b) KLD bound. By Theorem 1, given a small $\varepsilon$, for the budget $m:=C^{*}(\varepsilon)$ there exists $T\in\mathcal{F}_{m}$ such that $R\big(q_{T}\big)\;\leq\;\varepsilon.$ Because $\hat{q}_{m}$ minimizes $R(q)$ over $q\in\mathcal{Q}_{m}$,
@@
Since $\mathcal{Q}_{m}=\{q_{T_{\vartheta}}:\vartheta\in\Theta_{m}\}$, with a slight abuse of notation, we can write $R(q)$ as $R(\vartheta)$ and $R_{N}(q)$ as $R_{N}(\vartheta)$ for the corresponding $\vartheta$. As $\Theta_{m}$ is compact and $\vartheta\mapsto R_{N}(q_{\vartheta})$ is continuous, once again by Weierstrass Extreme Value Theorem, we have $\hat{q}_{N,m}\in\arg\min_{\mathcal{Q}_{m}}R_{N}(q)$ exists.
@@
Let $p(\theta\,|\,\eta)$ satisfies Assumption 1 for some compact $K_{\eta}$ and $p(\eta)$ is a distribution supported on $K_{\eta}$. Let $\{\eta_{r}\}_{r\geq 1}$ be a stationary ergodic Markov chain with invariant distribution $p(\eta)$. Let $\mathcal{F}_{m}=\{T_{\vartheta}|\;\vartheta\in\Theta_{m}\}$ denote a class satisfying Assumption 2. Let $q_{\vartheta}$ to be the conditional distribution $\theta\,|\,\eta=T_{\vartheta}(\eta,Z)$, $Z\sim N(0,1)$, and
$f_{\vartheta}(\eta)=\mathrm{KL}\big(q_{\vartheta}(\theta\mid\eta)\,\|\,p(\theta\mid\eta)\big),\qquad R(\vartheta)=\mathbb{E}[f_{\vartheta}(\eta)],\quad R_{N}(\vartheta)=\frac{1}{N}\sum_{r=1}^{N}f_{\vartheta}(\eta_{r}).$
@@
Envelope. Let $H(\vartheta,\eta,z)=\log\phi(z)-\log p\big(T_{\vartheta}(\eta,z)\mid\eta\big)-\log\partial_{z}T_{\vartheta}(\eta,z)$. Then $f_{\vartheta}(\eta)=\mathbb{E}_{\theta\,|\,\eta=T_{\vartheta}(\eta,Z)}[H(\vartheta,\eta,Z)]$. Following the proof of Corollary 1, $f_{\vartheta}(\eta)\leq\tau$ for some positive $\tau$ free of $\eta$ and for each fixed $\eta$, the map $\vartheta\mapsto f_{\vartheta}(\eta)$ and is continuous on compact $\Theta_{m}$, implying that it is uniformly continuous. Define for $\varepsilon>0$
@@
Then by uniform continuity $L_{\varepsilon}(\eta)\to 0$ pointwise as $\varepsilon\downarrow 0$, and $0\leq L_{\varepsilon}(\eta)\leq 2\tau$.
@@
$\mathbb{E}\,L_{\varepsilon}(\eta)\longrightarrow 0\qquad(\varepsilon\downarrow 0).$ (S8)
Since $\{\eta_{r}\}_{r\geq 1}$ is stationary ergodic and $L_{\varepsilon}\in L^{1}(p)$, Birkhoff’s ergodic theorem yields
$\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta_{r})\;\xrightarrow[N\to\infty]{a.s.}\;\mathbb{E}_{p(\eta)}[L_{\varepsilon}(\eta)].$
@@
$\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\big|R_{N}(\vartheta)-R_{N}(\vartheta_{j})\big|+\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\big|R(\vartheta)-R(\vartheta_{j})\big|.$
$\leq\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta_{r})+\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\mathbb{E}\,L_{\varepsilon}(\eta).$
@@
$\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\frac{1}{N}\sum_{r=1}^{N}L_{\varepsilon}(\eta_{r})+\max_{1\leq j\leq n(\varepsilon)}\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+\mathbb{E}\,L_{\varepsilon}(\eta).$
Taking $\limsup_{N\to\infty}$ on both sides and applying the ergodic theorem to the empirical average of $L_{\varepsilon}(\eta_{r})$ yields
$\limsup_{N\to\infty}\sup_{\vartheta\in\Theta_{m}}\big|R_{N}(\vartheta)-R(\vartheta)\big|$ $\leq\max_{1\leq j\leq n(\varepsilon)}\limsup_{N\to\infty}\big|R_{N}(\vartheta_{j})-R(\vartheta_{j})\big|+2\,\mathbb{E}\,L_{\varepsilon}(\eta).$ (S9)
For each fixed $j$, the ergodic theorem gives $R_{N}(\vartheta_{j})\to R(\vartheta_{j})$ almost surely, hence the first term on the right side of (S9) is zero. Therefore
@@
Let $\hat{q}_{m}\in\arg\min_{\mathcal{Q}_{m}}R(q)$, which exists by Corollary 1. Then, for all $M$,
$R(\hat{q}_{N,m})$ $\leq R_{N}(\hat{q}_{N,m})+\Delta_{N,m}\leq R_{N}(\hat{q}_{m})+\Delta_{N,m}\leq R(\hat{q}_{m})+2\Delta_{N,m}.$